You purchase a home and need to borrow $350,000. The bank is offering a 30-year loan that requires monthly payments and has a stated interest rate of 9% per year. What is your monthly mortgage payment? Now suppose that you can only afford to pay $2,500 per month. The bank agrees to allow you to pay this amount each month, yet still borrow the original amount. At the end of the mortgage in 30 years, you must make a balloon payment, that is, you must pay the remaining balance on the mortgage valued at year 30. What is the amount of the balloon payment you must make at the end of the mortgage? Provide the actual dollar amount of the balloon payment at year 30.
Answer #1

Part 1) Find the payment when there is no constraint on the amount of payment:

We are given the following information:

r 9.00%
n 30
frequency 12 (monthly payments)
PV $   3,50,000.00

We need to solve the following equation to arrive at the required monthly payment (PMT):
1-(1+ Frequency --)-nx frequency PV = PMT X- frequency 1-(1 + 0,09) -30x12 PV = PMT X- PMT = 2816.18
So the monthly payment is $2816.18

Part 2) Now this will require two part solution

  1. First we will have to find the FV of the loan amount at the end of 30 years
  2. Find the FV of the 2500 monthly payment at the end of year 30
  3. Subtract the FV of monthly payment from the FV of the loan amount
  4. The resultant amount is balloon payment

Step 1) We are given the following information:

PV $        3,50,000.00
r 9.00%
n 30
frequency 12

We need to solve the following equation to arrive at the required FV

FV = PV X (1+1 frequency frequencyxn 0.09 12x30 FV = 350000 X (1+. 12 FV = 5, 155, 701.64

So the FV is $5155701.64

Step 2) We are given the following information:

PMT $              2,500.00
r 9.00%
n 30
frequency 12

We need to solve the following equation to arrive at the required FV
(1+ Frequency) T ynx frequency - 1 FV = PMT X- frequency (1 + (1 + 0.09 30x12 - 1 12) FV = PMTX FV = 4576858.71

So the FV is $4576858.71

Step 3) Difference in the amount of two FVs is 5155701.64 - 4576858.71 =  $578,842.94

Step 4) Balloon payment =  $578,842.94

Below is the amortization Schedule:

Month Opening Balance PMT 2 3,50,125.00$ 2,500.00 | 3 و 350 , 250 . 943 2 , 500 . 00 $ 5$ 3,50,505.65 | 2,500.00$ واه و7 3 ,

  • Opening balance = previous year's closing balance
  • Closing balance = Opening balance+Loan-Principal repayment
  • PMT is calculated as per the above formula
  • Interest = 0.09 /12 x opening balance
  • Principal repayment = PMT - Interest
  • As the interest is higher than the Payment amount so the principal repayment will go increasing which needs to be paid off at the end to the last month in the form of a Balloon payment
  • Balloon payment is 578842.94 as calculated earlier.